OpenStudy (anonymous):

A train 120 metres in length passes a pole in 4 seconds and another train of length 210 metres travelling in the same direction in 15 seconds. What is the speed of the second train?

OpenStudy (anonymous):

First fine the speed of the first train

OpenStudy (anonymous):

It would be the distance traveled by the train divided by the time it took.

OpenStudy (anonymous):

For the second part, remember the second train has some speed \(v\). The first train will have to travel not only \(210\), but also \(120\) for it's own length and \(15\) for the distance the second train travels.

OpenStudy (anonymous):

\(15 \times v\) I mean.

OpenStudy (anonymous):

If the first train had a speed of \(w\), I would say: \[ w\times 4 = 120\\ w\times 15 = 210+120+15\times v \]

OpenStudy (anonymous):

first train speed =120/4=30m/s @wio i have a doubt that second train also 210/15 =70/5=14m/s

OpenStudy (anonymous):

confusing..........

OpenStudy (anonymous):

The first train passes the second train.

OpenStudy (anonymous):

you found that \[w=30\]

OpenStudy (anonymous):

Now use: \[ 30\times 15 = 210+120+15\times v \]to solve for \(v\)

OpenStudy (anonymous):

v=8m/s

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

what is meant by relative speed

OpenStudy (anonymous):

The relative speed is the difference in speed between the observer and the other moving object.